John invested $20,000 in a 14 percent annual interest yielding project. Deduce the compound amount of the project after 6 years if interest is compounded (a) quarterly, (b) weekly (c) thirdly (d) monthly (e) daily
Answer by Boreal(15235)

P=Po(1+r/n)^nt, where r is the rate, n the number of compoundings per year, and nt the total number of compoundings.
20000(1+0.14/4)^24, because 24 compoundings quarterly in 6 years. Do the raising of the power first and round at the end. $45666.57
-------------
weekly
20000*(1+(.14/52))^312=$46275.08
------------
thirdly
20000*(1+.14/3)^18=$45454.94
-----------
monthly
20000*(1+.14/12)^72=$46102.63
----------
daily
20000*(1+0.14/365))^2190=$46319.88